# LLM-Craft: Robotic Crafting of Elasto-Plastic Objects with Large Language Models

### v2

@@ Abstract
When humans create sculptures, we are able to reason about how geometrically we need to alter the clay state to reach our target goal. We are not computing point-wise similarity metrics, or reasoning about low-level positioning of our tools, but instead determining the higher-level changes that need to be made. In this work, we propose LLM-Craft, a novel pipeline that leverages large language models (LLMs) to iteratively reason about and generate deformation-based crafting action sequences. We simplify and couple the state and action representations to further encourage shape-based reasoning. To the best of our knowledge, LLM-Craft is the first system successfully leveraging LLMs for complex deformable object interactions. Through our experiments, we demonstrate that with the LLM-Craft framework, LLMs are able to successfully reason about the deformation behavior of elasto-plastic objects. Furthermore, we find that LLM-Craft is able to successfully create a set of simple letter shapes. Finally, we explore extending the framework to reaching more ambiguous semantic goals, such as ”thinner” or ”bumpy”. For videos and prompting details, please visit our project website: https://sites.google.com/andrew.cmu.edu/llmcraft/home.
@@ I Introduction
With the goal of building robust and generalizable autonomous robotic systems, we need to design frameworks that can reason about complex interactions between the robot and the environment. Deformable object manipulation is a great test task for reasoning about these complex interactions, as it requires understanding how the state of the object itself will change during contact. In this work, we will investigate the task of deformation-based crafting, in which the robot moulds the clay into a variety of simple shapes with a parallel gripper. This task allows us to investigate the challenges of both predicting the behavior of the clay as well as long-horizon planning and reasoning. Previous work focused on elasto-plastic object manipulation and crafting typically work at the very low-level, predicting clay dynamics [shi2022robocraft, bartsch2023sculptbot, shi2023robocook], or direct action imitation [bartsch2024sculptdiff]. However, these approaches do not leverage existing world knowledge to reason about how clay behaves, or have the ability to reason at a geometric or semantic level.
@@
In this work, we present LLM-Craft, a preliminary study exploring how large language models (LLMs) reason about the complex task of elasto-plastic object crafting. LLMs have been demonstrated to contain useful world knowledge for a wide range of reasoning and generation tasks [mirchandani2023large, wu2023smartplay, todd2023level, yu2023language, jadhav2024large, wang2024can]. We hypothesize that LLMs have substantial and useful world knowledge for the crafting task, such as understanding how clay will behave when pressed, and understanding semantic shapes and goals. To better facilitate higher-level geometric reasoning for crafting action selection, we simplify our state and action representation substantially. Through extensive real-world experiments, we demonstrate the success of leveraging LLMs directly for the clay crafting task. We find that the ability of LLMs to reason at the geometric and semantic level is particularly powerful and presents an argument for future crafting pipelines incorporating LLMs. The key contributions of this work are as follows.
  • To the best of our knowledge, we present the first system successfully using LLMs for the real-world elasto-plastic object crafting task.
  • We conduct an exploration of the semantic reasoning capabilities of LLMs and the usefulness of this higher-level reasoning for crafting.
  • We demonstrate that with careful prompt engineering, LLMs can in fact successfully reason about complex robot-object interactions.
Figure 1: INSERT CAPTION.
@@
Figure 2: Method Pipeline. The LLM-Craft system takes a top-down image of the clay with a wrist-mounted camera as the state observation. A grid is applied to the image to represent the discrete regions of the clay for the LLM to reason about where to grasp. The LLM is then prompted with the gridded state and goal images as well as an action prompt. The LLM selects a sequence of grasps to apply to the clay, and the robot executes the first one. A new state observation is collected, and passed back to the LLM along with the goal image and the termination prompt to determine if the goal has been reached. If the goal has not been reached, the LLM is then queried with the state and goal images as well as the action prompt. This is an iterative process until the goal has been reached.
@@ II Related works
Elasto-Plastic Object Manipulation: There have been a variety of successful past works within the realm of elasto-plastic object manipulation and crafting. There are numerous simulations modeling the deformation behavior of these objects, but there remains a large sim-to-real gap [heiden2021disect, hu2018moving, huang2021plasticinelab, gu2023maniskill2]. These simulations enabled trajectory optimization-based approaches that were able to bridge the sim-to-real gap for simple tasks [qi2022learning, yamada2024d]. Many real-world clay crafting frameworks learn the low-level dynamics of the clay to be used to generate action trajectories [shi2022robocraft, shi2023robocook, bartsch2023sculptbot, bauer2024doughnet, li2024deformnet]. Alternatively, a range of imitation-based frameworks have been successful, where demonstrations are abstracted into skills [lin2022planning, li2023dexdeform], or more direct trajectory imitation [bartsch2024sculptdiff]. All of these successful real-world approaches require hardware and task-specific datasets, with many requiring human demonstrations which are particularly time consuming to collect. Additionally, these methods are very brittle to particular design choices and are not always transferable across lab setups. In this work, we investigate if we can leverage the general world knowledge of LLMs to shape elasto-plastic objects with a parallel gripper. From the lens of the deformable manipulation community, if LLMs do indeed have useful world knowledge to make high-level predictions of the objects’ behavior, this could be a useful component to future pipelines that could improve generalizability.
@@
LLMs for Robotic Tasks: There is a significant body of previous work leveraging LLMs for a variety of robotic tasks. LLMs have been particularly effective as robotic task planning agents [wu2024mldt, wang2023describe, bhat2024grounding, rana2023sayplan, dorbala2023can, sharan2024plan, singh2023progprompt, ding2023task, gupta2024action, cheng2024empowering] with many recent works finding success incorporating a variety of real-world grounding techniques including scene graphs [rana2023sayplan], diffusion models [sharan2024plan], counterfactual perturbation [wang2024grounding], developing a multimodal language model for embodied tasks [driess2023palm] and 3D value maps [huang2023voxposer]. These LLM-based frameworks span a wide range of reasoning levels, from low-level autoregressive code policy generation [liang2023code] to reasoning about task plans purely with natural language [mikami2024natural]. While the widespread success of LLM robotic methods has been impressive, most research has focused on the relatively simple tasks of navigation [dorbala2023can, shah2023lm, rana2023sayplan, yang2023llm], or sequences of pick and place actions [singh2023progprompt, sharan2024plan, wu2024mldt, bhat2024grounding, gupta2024action]. These tasks do require long-horizon reasoning as well as recognition and interaction with a variety of real-world objects, but the general action skillset is relatively simple and does not require much intuition and understanding of physical robot-object interactions. In Tidybot [wu2023tidybot], researchers present an LLM planning framework that involves interacting with deformable cloth, but the specific sorting tasks explored do not require reasoning about the complex deformable dynamics and the robot-object interactions remain only pick and place. Similarly, in [xu2023creative] researchers explore how LLMs can creatively reason about tool use to achieve a set of tasks, but the tool-object interactions remain very simple. In fact, in [kwon2023language] researchers concluded that LLMs often simplify or incorrectly predict more complicated interactions between the robot and objects. However, the general motivation behind the significant body of LLM-based robotic methods is that they are leveraging the world knowledge that these LLMs contain to reason about the task. In this work, we hypothesize and demonstrate that LLMs do indeed have the ability to predict and reason about complex robot-object interactions with the task of the robotic crafting. We particularly explore how LLMs can plan complicated action sequences that require reasoning about the effect of the grasp action on the objects’ state itself.
@@ III Methodology
The key components of LLM-Craft are the choice of state and action representations as well as the iterative prompting scheme with the LLM. An overview of the pipeline is shown in Figure 2. In this work, we are using Gemini 1.0 Pro Vision [team2023gemini] as our multimodal large language model that can take both images and text as input. We chose this version of Gemini over the more recent Gemini 1.5 [reid2024gemini] models as Gemini 1.0 Pro Vision was optimized for generating image descriptions and identifying objects in images which is a key part of the reasoning task for crafting shapes. Additionally with preliminary experiments we found it to substantially outperform Gemini 1.5. We have not investigated alternative LLMs, but the LLM-Craft pipeline can be used with any LLM that can take both images and text as input, such as GPT-4 [achiam2023gpt].
@@ III-A State and Action Representation
In this work we are proposing an LLM-based framework for the task of top-down shape crafting of elasto-plastic objects with a parallel gripper, meaning the only actions the system is able to apply to the clay is a squeeze of a fixed distance between the fingertips. In existing robotic clay crafting literature, the common choice of state representation is a point cloud, and the action representation is the continuous end-effector pose [shi2022robocraft, bartsch2023sculptbot, shi2023robocook, bartsch2024sculptdiff]. While these state and action choices are very effective in existing pipelines, they can disconnect the reasoning about actions from the state representation. When reasoning about how to shape clay, humans determine at a high level where and how the clay needs to be deformed. In this work, we propose a simple gridded top-down image of the clay as the state and action space to allow the LLM to reason about the larger geometric regions of the clay without needing to directly reason about global end effector poses.
@@
The gridded state space is visualized in Figure 2. We equip the robot with an end-effector mounted camera to capture top-down images of the clay. We then superimpose a simple 4x4 grid to the image to break the clay up into different geometrical regions. Each cell in the grid is 2cm by 2cm in the robot frame, as the 3D printed fingertips on the robot are approximately 2cm in width. A visualization of the hardware setup is shown in Figure 3. We represent the columns of the grid from left to right as ’A’, ’B’, ’C’, and ’D’ and the rows from top to bottom as ’1’, ’2’, ’3’, and ’4’, thus cells individually are referenced as ’A1, ’B1’, etcetera. The action space is defined as selecting two cells to push a fixed distance towards each other. This choice of action representation significantly simplifies action selection as there is a discrete set of actions for the LLM to choose from, which is a much simpler task than selecting continuous action parameters. Additionally, this action representation is more closely linked to predicting the clay deformation behavior, than directly selecting the global end-effector pose. Through experiments we find that while this representation drastically constrains the action space, both the LLM-Craft and a human baseline are able to successfully create a variety of shape goals. At execution time, we assume the clay position remains fixed on the elevated stage, and thus know the global position of each of the grid points to transform the discrete cells selected by the LLM into an end-effector pose for squeezing.
III-B Prompting
The goal of this work is to explore how we can leverage the world knowledge and shape-level reasoning capabilities of LLMs without requiring any further training or fine-tuning on our specific task. The full process diagram of the iterative prompting scheme is shown in Figure 2, in which we use an action prompt and a termination prompt. The role of the action prompt is to query the LLM with an explanation of the task, state and action space, and the robot embodiment to then output a crafting action sequence of varying length. We take inspiration from past successful approaches using LLMs for robotic tasks for some of the prompting details [kwon2023language]. The components of the action prompt are as follows: (1) an overview explaining the robot embodiment and the high-level details of the crafting task, (2) an explanation of the environmental setup including the clay and the grid coordinate system, (3) an explanation of the grasp action, including that the end-effector pose cannot change during the grasp, (4) an explanation of how to generate a crafting trajectory, including the requirements of (a) describing the current and goal shapes, (b) reasoning about the grasp strategy, (c) describe the predicted effect of the chosen grasp to the clay state, (5) the goal gridded image, (6) the current state gridded image, and (7) the user command to create the goal from the current state. The role of the termination prompt is to query the LLM to determine if the goal shape has been reached. The advantage of using an LLM as opposed to similarity metric thresholding is that we are prompting the LLM to judge the shape similarity more holistically, i.e. if both of the shapes are of the letter T, which is a fundamentally different task from judging point-wise similarity between point clouds. The terimation prompt consists of (1) an overview of the shape comparison task, (2) an explanation of how to generate a decision, with the requirements to (a) describe the similarities and differences between the state and goal, and (b) make a decision if crafting should continue, (3) the current gridded state image, (4) the goal gridded image, and (5) the user command to determine if crafting should stop. For full details of both prompts, please see the Supplemental Materials. In section IV-E we will discuss prompt ablations to evaluate the impactfullness of each component of the prompts on the final system performance.
@@ IV Experiments
To fully evaluate LLM-Craft we conduct a variety of experiments, particularly with single-step shape goals and longer horizon letter shape goals evaluating the system performance as well as the impacts of the choice of state representation and the ability for the system to replan. Additionally, we conduct an ablation study to evaluate the importance of each component of the prompt on the final performance. Given the results of our experiments, we further explore how LLM-Craft can be extended to handle more semantic goals. For each experiment, we conduct five real-world runs per goal and report the mean and standard deviation of the Chamfer Distance (CD) and Earth Mover’s Distance (EMD) between the point clouds of the final shape and the goal. The hardware setup is shown in Figure 3. With a wrist-mounted RGB-D camera, we record the top-down image of the clay before and after each grasp. Additionally after the crafting sequence is complete, we record a top-down point cloud of the clay for our quantitative similarity metric evaluations compared to a point cloud of the goal shape.
IV-A State Representation
The choice of state representation is a gridded top-down image of the clay. However, to explore alternative state representations that are compatible with the grid-based action space, we implemented a baseline in which we represent the clay state as a binary array with a 1 if more than half of the cell contains clay and with a 0 if less than half the cell contains clay. While both state representations provide sufficient information for the crafting task, we argue that the image-based respresentation allows for shape-level reasoning more easily. To determine the advantages/disadvantages of these state representations, we conducted a set of single-step experiments in which the LLM needed to select a single grasp to reach the goal state from the current state. While both the gridded image and binary array representations have similar CD and EMD metrics with respect to the goal, visually the gridded image representation is able to more consistently match the goal shape. While the LLM can get confused about rotation with both representations, this happens less frequently with the gridded image. The image representation inherently has much more information than the simple binary array.

### v3

@@ Abstract
When humans create sculptures, we are able to reason about how geometrically we need to alter the clay state to reach our target goal. We are not computing point-wise similarity metrics, or reasoning about low-level positioning of our tools, but instead determining the higher-level changes that need to be made. In this work, we propose LLM-Craft, a novel pipeline that leverages large language models (LLMs) to iteratively reason about and generate deformation-based crafting action sequences. We simplify and couple the state and action representations to further encourage shape-based reasoning. To the best of our knowledge, LLM-Craft is the first system successfully leveraging LLMs for complex deformable object interactions. Through our experiments, we demonstrate that with the LLM-Craft framework, LLMs are able to successfully create a set of simple letter shapes. We explore a variety of rollout strategies, and compare performances of LLM-Craft variants with and without an explicit goal shape images. For videos and prompting details, please visit our project website: https://sites.google.com/andrew.cmu.edu/llmcraft/home.
@@ I Introduction
With the goal of building robust and generalizable autonomous robotic systems, we need to design frameworks that can reason about complex interactions between the robot and the environment. Deformable object manipulation is a great test task for reasoning about these complex interactions, as it requires understanding how the state of the object itself will change during contact. In this work, we will investigate the task of deformation-based crafting, in which the robot molds the clay into a variety of simple shapes with a parallel gripper. This task allows us to investigate the challenges of both predicting the behavior of the clay as well as long-horizon planning and reasoning. Previous work focused on elasto-plastic object shaping typically work at the very low-level, predicting clay dynamics [1, 2], leveraging human-created sub-goals to speed up action planning [3] or direct action imitation [4]. However, these approaches do not leverage existing world knowledge to reason about how clay behaves or have the ability to reason at a geometric or semantic level. Furthermore, existing methods for the clay shaping task rely heavily on hardware-specific real-world datasets which can be very time consuming to collect and are often not transferable across hardware configurations [1, 3, 2, 4, 5].
@@
In this work, we aim to explore to what extent we are able to directly leverage Large Language Models (LLMs) without any task-specific data or finetuning for this highly complex manipulation task. Particularly, do LLMs contain useful world knowledge to generate long-horizon action sequences to create simple letter shapes with deformation-based actions? LLMs have been demonstrated to contain useful world knowledge for a wide range of reasoning and generation tasks [6, 7, 8, 9, 10, 11]. However, past work using LLMs to generate plans for robotic tasks typically focus on simple pick and place actions that require less sophisticated prediction of future states [12, 13, 14, 15, 16]. We present LLM-Craft, a preliminary study exploring how LLMs reason about the complex task of elasto-plastic object crafting. We hypothesize that LLMs have substantial and useful world knowledge to generate action sequences for the clay crafting task, such as understanding how clay will behave when pressed or squeezed and understanding simple semantic shapes and goals. To better facilitate higher-level geometric reasoning for crafting action selection, we simplify our state and action representation substantially. Through extensive real-world experiments, we demonstrate the success of leveraging LLMs directly for the clay crafting task. We find that the ability of LLMs to reason at the geometric and semantic level is particularly powerful and presents an argument for future crafting pipelines incorporating LLMs.
@@
Figure 1: Method Pipeline. a) The LLM-Craft system takes a top-down image of the clay with a wrist-mounted camera as the state observation. A grid is applied to the image to represent the discrete regions of the clay. The LLM is prompted with the gridded state/goal images and an action prompt. The LLM selects a sequence of grasps to apply to the clay, and the robot executes the first one. A new state observation is collected and passed back to the LLM along with the goal image and the termination propmt to determine if the goal has been reached. If the goal has not been reached, the LLM is iteratively re-queried until the goal has been achieved. b) A visualization of different method deployment strategies. We present chain of thought [17] as our iterative base method, and self-consistency (SC) [18] and tree of thought (ToT) [19] are rollout strategies that require multiple queries of the LLM per action trajectory generation.
@@ II Related works
Elasto-Plastic Object Manipulation: There have been a variety of successful past works within the realm of elasto-plastic object manipulation and crafting. There are numerous simulations modeling the deformation behavior of these objects, but there remains a large sim-to-real gap [20, 21, 22, 23]. These simulations enabled trajectory optimization-based approaches that were able to bridge the sim-to-real gap for simple tasks [24, 25]. Many real-world clay crafting frameworks learn the low-level dynamics of the clay to be used to generate action trajectories [1, 3, 2, 5, 26]. Alternatively, a range of imitation-based frameworks have been successful, where demonstrations are abstracted into skills [27, 28], or more direct trajectory imitation [4]. All of these successful real-world approaches require hardware and task-specific datasets, with many requiring human demonstrations which are particularly time consuming to collect. Additionally, these methods are very brittle to particular design choices and are not always transferable across lab setups. In this work, we investigate if we can leverage the general world knowledge of LLMs to shape elasto-plastic objects with a parallel gripper. From the lens of the deformable manipulation community, if LLMs do indeed have useful world knowledge to make high-level predictions of the objects’ behavior, this could be a useful component to future pipelines that could improve generalizability.
@@
LLMs for Robotic Tasks: There is a significant body of previous work leveraging LLMs for a variety of robotic tasks. LLMs have been particularly effective as robotic task planning agents [14, 29, 15, 30, 31, 13, 12, 32, 16, 33, 34] with many recent works finding success incorporating a variety of real-world grounding techniques including scene graphs [30], diffusion models [13], counterfactual perturbation [35], developing a multimodal language model for embodied tasks [36] and 3D value maps [37]. These LLM-based frameworks span a wide range of reasoning levels, from low-level autoregressive code policy generation [38] to reasoning about task plans purely with natural language [39]. While the widespread success of LLM robotic methods has been impressive, most research has focused on the relatively simple tasks of navigation [31, 40, 30, 41], or sequences of pick and place actions [12, 13, 14, 15, 16]. These tasks do require long-horizon reasoning as well as recognition and interaction with a variety of real-world objects, but the general action skillset is relatively simple and does not require much intuition and understanding of physical robot-object interactions. In Tidybot [42], researchers present an LLM planning framework that involves interacting with deformable cloth, but the specific sorting tasks explored do not require reasoning about the complex deformable dynamics and the robot-object interactions remain only pick and place. Similarly, in [43] researchers explore how LLMs can creatively reason about tool use to achieve a set of tasks, but the tool-object interactions remain very simple. In fact, in [44] researchers concluded that LLMs often simplify or incorrectly predict more complicated interactions between the robot and objects. In this work, we hypothesize that LLMs contain useful world knowledge to generate coherent sculpting sequences, such as needing to conduct sequential perpendicular grasps to create an X shape. In this work, we explore how LLMs can plan action sequences involving complex robot-object interactions by predicting an action sequence as well as the predicted effect of the grasp action on the objects’ state itself.
@@ III Methodology
The key components of LLM-Craft are the choice of state and action representations as well as the iterative prompting scheme with the LLM. An overview of the pipeline is shown in Figure 1. In this work, we are using Gemini 1.0 Pro Vision [45] and Gemini 2.0 Flash [46] as our multimodal large language models that can take both images and text as input. Due to the fast-moving nature of state-of-the-art LLMs, model versions were depreciated through our experiment revisions. We have done our best to allow for comparison between model versions and experiment ablations. We have not investigated alternative LLMs, but the LLM-Craft pipeline can be used with any LLM that can take both images and text as input, such as GPT-4 [47].
@@ III-A State and Action Representation
In this work we are proposing an LLM-based framework for the task of top-down shape crafting of elasto-plastic objects with a parallel gripper. In existing robotic clay crafting literature, the common choice of state representation is a point cloud, and the action representation is the continuous end-effector pose [1, 2, 3, 4]. While these state and action choices are very effective in existing pipelines, they can disconnect the reasoning about actions from the state representation. When reasoning about how to shape clay, humans determine at a high level where and how the clay needs to be deformed. In this work, we propose a simple gridded top-down image of the clay as the state and action space to allow the LLM to reason about the larger geometric regions of the clay without needing to directly control global end-effector poses. The gridded state space is visualized in Figure 1. We equip the robot with an end-effector mounted camera to capture top-down images of the clay. We then superimpose a simple 4x4 grid onto the image to break the clay into different geometrical regions. Each cell in the grid is 2cm by 2cm in the robot frame, as the 3D printed fingertips on the robot are 2cm in width. A visualization of the hardware setup is shown in Figure 2. We represent the columns of the grid from left to right as ’A’, ’B’, ’C’, and ’D’ and the rows from top to bottom as ’1’, ’2’, ’3’, and ’4’, thus cells individually are referenced as ’A1, ’B1’, etcetera. The action space is defined as selecting two cells to push a distance towards each other. This choice of action representation significantly simplifies action selection as there is a discrete set of actions for the LLM to choose from, which is a much simpler task than selecting continuous action parameters. Additionally, this action representation is more closely linked to predicting the clay deformation behavior, than directly selecting the global end-effector pose. Through experiments we find that while this representation drastically constrains the action space, both the LLM-Craft and a human baseline are able to successfully create a variety of shape goals. At execution time, we assume the clay position remains fixed on the elevated stage, and thus know the global position of each of the grid points to transform the discrete cells selected by the LLM into an end-effector pose for squeezing.
III-B Prompting
In this work, we explore how we can leverage the world knowledge and shape-level reasoning capabilities of LLMs without requiring any data collection, further training or fine-tuning on our specific task. The full process diagram of the iterative prompting scheme is shown in Figure 1, in which we use an action prompt and a termination prompt. The action prompt provides the LLM with an explanation of the task, state/action space, and the robot embodiment, and requires the LLM to output a crafting action sequence of varying length. We take inspiration from past successful approaches using LLMs for robotic tasks for some of the prompting details [44]. The components of the action prompt are as follows: (1) an overview of the robot embodiment and the high-level details of the crafting task, (2) the environmental setup including the clay and the grid coordinate system, (3) the grasp action, including that the end-effector pose cannot change during the grasp, (4) how to generate a crafting trajectory, including the requirements of (a) describing the current and goal shapes, (b) reasoning about the grasp strategy, (c) describe the predicted effect of the chosen grasp to the clay state, (5) the goal gridded image, (6) the current state gridded image, and (7) the user command to create the goal from the current state. The role of the termination prompt is to query the LLM to determine if the goal shape has been reached. The termination prompt consists of (1) an overview of the shape comparison task, (2) how to generate a decision, with the requirements to (a) describe the similarities and differences between the state and goal, and (b) make a decision if crafting should continue, (3) the current gridded state image, (4) the goal gridded image, and (5) the user command to determine if crafting should stop. For full details of both prompts, please see the Supplemental Materials on the project website. In section IV-E we will discuss prompt ablations to evaluate the impactfullness of each component of the prompts on the final system performance.
III-C Rollout Strategies
In our experiments, we explore standard iterative querying of the LLM with Chain of Thought prompting [17], as well as more sophisticated strategies to mitigate error, such as Self-Consistency (SC) [18] and Tree of Thought (ToT) [19]. For our Self-Consistency implementation, we leverage sampling-based generation in which we query the LLM 10 times to generate candidate sculpting trajectories, and execute the most frequently generated trajectory. Due to the symmetrical nature of our squeezing actions, we consider ’squeeze A1 and B2’ and ’squeeze B2 and A1’ to be the same actions. For our breadth-first-search Tree of Thought implementation, we leverage proposal-based generation in which we prompt the LLM to propose 4 possible actions given the current state and action history. This better ensures that we generate a wide diversity of candidate sculpting trajectories without generating any duplicates. Additionally, we employ voting-based value prediction of each candidate trajectory in the tree, as it is very difficult to quantify the value of candidate sculpting trajectories directly.
@@ IV Experiments
To fully evaluate LLM-Craft we conduct a variety of experiments, particularly with single-step shape goals and longer horizon letter shape goals evaluating the system performance as well as the impacts of the choice of state representation, grid size, image versus text-based goals, and the rollout strategy. Additionally, we conduct an ablation study to evaluate the importance of each component of the prompt on the final performance, as well as the impact of varying grid sizes. For each experiment, we conduct five real-world runs per goal and report the mean and standard deviation of all quantitative metrics. The hardware setup is shown in Figure 2. With a wrist-mounted RGB-D camera, we record the top-down image of the clay before and after each grasp. Additionally after the crafting sequence is complete, we record a top-down point cloud of the clay for our quantitative similarity metric evaluations compared to a point cloud of the goal shape.
IV-A Quantitative Evaluation Metrics
Chamfer Distance (CD) and Earth Mover’s Distance (EMD) between the final point cloud and the goal shape are the most common quantitative metrics used to evaluate the clay shaping task in literature [3, 1, 2, 4]. However, for the more semantic task of shaping a specific letter, CD and EMD do not fully capture the nuanced qualities humans associate with a ’good’ clay shape. Therefore, in addition to reporting CD/EMD, we propose additional metrics to represent the curvature smoothness of the shapes created. We represent the ”smoothness” of the created shapes with the equation of curvature of parametric curves (equation 1), and with the perimeter-to-area ratio (PAR) on the 2D outline of each shape extracted with Canny edge detection.
$curvature=\frac{||x^{\prime\prime}y^{\prime}-y^{\prime\prime}x^{\prime}||}{(x^{\prime 2}+y^{\prime 2})^{\frac{3}{2}}}$ (1)
Beyond these quantitative metrics, we conducted a human evaluation study following the procedure of RoboCook [3] to further evaluate the performance of LLM-Craft. To evaluate all 11 methods (in II) over the 5 different shapes (C, I, L, T, X), we created 11 randomized surveys each with a single instance of each letter and a random method. This ensured that each survey taker only saw a single instance of each letter, as to not bias their responses. We asked each respondent to first classify the letter in the provided image, and then rate the shape quality on a scale from 1-10.